OpenStudy (anonymous):

the slope of the line x+9=y is

OpenStudy (anonymous):

1

OpenStudy (anonymous):

this equation is in the form y=mx+b. m is the slope and the coefficient of x in your equation is 1

OpenStudy (anonymous):

sorry but im not sure what that means

OpenStudy (anonymous):

from what corey said, y=mx+b. that is the standard equation of a line. where (from left to right) y=output m=slope x= x input b= vertical intercept so you could have y=4x+3 in which case 4=m and 3=b and 4 would be your slope for that equation. and to relate it back to your original problem, y=x+9, there is no number in front of x, so it assumed to be 1. because 1x+9 is the same as x+9. and the y intercept (b) is simply 0,0

OpenStudy (anonymous):

nice elaboration